UN (aq) + H2O(1) = NO (aq) + OH(aq) Kb = 1.0 x 10-5 @ 25 °C H Benadryl BH a. How would the following actions disturb the equiThe answer to part b is pH= 10.70, %ionization= 2.0%.

The answer to part c is pH=9.22

Please show all work
Answer #1

LON- (ar) + 1,0 () = Danke # (Bit) ton Kb=100x105 The reaction is in equilibrium. According to Le chatelier principle when anAction AG lo recent Benachrig (BHB COH) CHA | PH ] -ward) Dissolving and negative Toward Increases Increases Increases decreaIf x is degree of dissociation then 8 + H2O = 8H + OH (1-age ac xe at eam. Kb = [BH], [or ] x c.ac (1-2).C Since, & LL1 So, 1O Mixture of Benadryl and ammonium Chloride Salt copie de buffer solution- hasselbalch equation From Henderson We can write-